OpenStudy (anonymous):

Find two positive real numbers whose product is a maximum. The sum of the first and three times the second is 30

OpenStudy (anonymous):

x + 3y = 30 y = 10 - x/3 Maximize f(x) = x(10 - x/3) = 10x - (x^2)/3 f'(x) = (d/dx)(10x - (x^2)/3) = 10 - 2x/3 Maximum at f'(x) = 0 10 - 2x/3 = 0 x = 15 y = 10 - 15/3 = 5

OpenStudy (anonymous):

What do you mean?